Circular No. 8139

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 K4 (LINEAR) An apparently asteroidal object found by the LINEAR survey (discov-

ery observation below), posted on the NEO Confirmation Page, has been found to show a round coma of diameter 5′′–7′′ (m_{1} = 17.5) on CCD images taken by J. Young on May 29.5 and 30.4 UT with the 0.6-m reflector at Table Mountain. J. McGaha, Tucson, AZ, reports that three stacked, 2-min CCD images, taken on May 29.4 with a 0.30-m reflector, show a 3′′ nuclear condensation and a 6′′ coma that is offset to the northeast.

2003 UT  α_{2000}  δ_{2000}  m_{2}
May 28.38281  21^{h}15^{m}22.^{s}08  +28^{o}54′27′′.0  17.6

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-K52.

T = 2003 Sept.11.935 TT  ω = 51.^{o}615
Ω = 344.449 }2000.0
q = 8.47456 AU  i = 111.645

COMET P/2003 KV_{2} (LINEAR) Another apparently asteroidal LINEAR object (discovery observation

below), announced on MPEC 2003-K27 as 2003 KV_{2} (see also MPEC 2003-K38 and 2003-K47), has been found cometary on R-band images taken by C. Brinkworth and M. Burleigh on May 28.9 and 29.9 UT with the 1-m Jacobus Kapteyn Telescope on La Palma (communicated by A. Fitzsimmons), in which the object shows a tail ∼ 4′′-5′′ long in p.a. 125^{o} and a small coma that is somewhat larger than the surrounding field stars.

2003 UT  α_{2000}  δ_{2000}  m_{2}
May 23.16257  10^{h}52^{m}41.^{s}18  +37^{o}50′29′′.1  18.4

Additional astrometry and the following orbital elements appear on MPEC 2003-K53. This orbit shows a passage 0.55 AU from Jupiter in Jan. 2001, before which the perihelion distance was somewhat larger.

T = 2003 July 10.882 TT  ω = 188.^{o}789
e = 0.62667  Ω = 66.478 }2000.0
q = 1.06234 AU  i = 25.501
a = 2.84558 AU n^{o} = 0.205327 P = 4.80 years
2003 May 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green